please help me with this equation: graph the following equation, make a solution set table with at least 3 solutions, find the x intercept and y intercept and slope of -2x+y=2
Found 2 solutions by solver91311, lenny460:
Answer by solver91311(24713)


Graph:

Step 1. Pick a value for x. It can be anything you like, but 0, 1, or some other small integer usually works well and makes the arithmetic easier.

Step 2. Substitute that value in place of x in your equation. Do the arithmetic and determine the value of y that results.

Step 3. Take the value of x that you selected for step 1 and the value of y that you calculated in step 2 and form an ordered pair (x,y).

Step 4. Plot the ordered pair from Step 3 on your graph. Remember that the x value is the distance right or left along the horizontal axis and the y value is the distance up or down along the vertical axis.

Step 5. Repeat steps 1 through 4 with a different value for x.

Step 6. Draw a line across your graph that passes through the two points that you plotted.

Make a solution set table: Use the coordinates from the two ordered pairs created above to make two of the solutions for your table. Then repeat steps 1 through 4 to create the third solution.

Intercepts: If one of your solutions above has an x-coordinate of 0, then you have the y-intercept already. If not, choose 0 for an x value and repeat steps 1 through 4 above. Then substitute 0 for y and solve for x to get the x-intercept. Remember, intercepts are points so they have to be specified with an ordered pair. The x-intercept looks like (a,0) and the y-intercept looks like (0,b)

Slope: Solve the equation for y in terms of everything else. In this case, all you have to do is add 2x to both sides. When the equation looks like y = mx + b, then the coefficient on x gives the slope.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lenny460(1073)

Plot the graph: -2x + y = 2
Add 2x to each side of the equation.
-2x + y + 2x = 2 + 2x
Combine like terms:
-2x + 2x = 0
Therefore:
y = 2 + 2x or
y = 2x + 2


From the above we get the following coordinates:
(-4,-6)
(-2,-2)
(-1,0)
(0,2)
(2,6)


From the graph:
The y-intercept = 2
The x-intercept = -1
The Slope = rise/run = 8/4 = 2
The Slope = 2

I hope the above explanations will help you to plot the graph. Sorry I don't have the software to plot the graph here. You can always contact me at:


obuong3@aol.com